OpenStudy (anonymous):

A fish pond at the local park is a regular hexagon. a. Write a formula for the perimeter of the pond in terms of the length of a side. Explain your formula. b. Each side has a length of 7.5 feet. Find the perimeter of the pond. c. Suppose the designer of the pond wants to make another regular hexagonal pond with a perimeter of 57.6 feet. Find the length of one side of the pond. Explain your method.

OpenStudy (anonymous):

a. perimeter = 6 sides*a = 6a

OpenStudy (anonymous):

your qstns are so lengthy

OpenStudy (anonymous):

for part b , just substitute 7.5 into the formula for part c: 57.6 = 6*a thererfore, a = 57.6/6